# Superquantile-Gibbs Relaxation for Minima-selection in Bilevel Optimization

### v2

@@ 3.2 Common Dimension and Geometric Properties of $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$
For each $\theta\in\Theta$, the local PŁ condition implies that on $\mathcal{S}(\theta)$ the Hessian $\partial_{x}^{2}g(\theta,x)$ is positive definite in the normal directions and degenerate in the tangent directions (rebjock2023fast, Proposition 2.7). In particular, its rank $\texttt{r}(\theta)$ is constant over $x\in\mathcal{S}(\theta)$, and the dimension of the minimizer manifold can be written as $\dim\mathcal{S}(\theta)=\texttt{k}(\theta)=d-\texttt{r}(\theta)$. Moreover, the solution mapping $\theta\mapsto\mathcal{S}(\theta)$ is continuous (Section˜3.1), and $\partial_{x}^{2}g(\theta,x)$ is continuous in $(\theta,x)$. Thus, when $\theta$ is perturbed slightly, points in $\mathcal{S}(\theta)$ move continuously and the corresponding Hessians remain close, so no eigenvalue can cross zero and the rank $\texttt{r}(\theta)$ cannot change. Hence $\texttt{r}(\theta)$, and therefore $\texttt{k}(\theta)$, is locally constant in $\theta$ and we obtain a single intrinsic dimension shared by all manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$. We summarize this in the following lemma.
@@ 4.1 Approximation of $F_{\max}$ with $\tilde{F}_{\texttt{SQ}}$
Classical results such as (gray1973volume) give a pointwise asymptotic expansion of the volume of a small geodesic ball, showing that $\mathcal{M}(B_{\mathsf{g}}(p;r))/r^{\texttt{k}}\to c(p)\,$ as $r\to 0$ with a coefficient depending on the local curvature at $p$. For our purposes, this is not sufficient: we will see that in the proof of ˜4.2 we need a uniform, non-asymptotic lower bound that holds simultaneously for all centers $p\in\mathcal{S}(\theta)$ and all radii $0<r\leq r_{1}$, with $r_{1}$ depending only on the global bounds on sectional curvature and injectivity radius from Section˜3.2. Section˜4.1 provides exactly this quantitative version by turning the asymptotic comparison into an explicit estimate with controlled radius and constants.
@@
The proof of Section˜4.1 is given in Section˜C.1. Based on Section˜4.1, the Riemannian volume of a geodesic ball with radius $r$, $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$, scales as $r^{\texttt{k}}$. At a high level, the lemma relies on two geometric ingredients: (i) The injectivity radius bound $r_{0}>0$ guarantees that, around every $p\in\mathcal{S}$, we can introduce normal coordinates on the whole geodesic ball $\mathbb{B}_{\mathsf{g}}(p;r)$ for all $r\leq r_{0}$. (ii) The curvature bound $|\mathrm{Sec}|\leq\mathsf{C}_{0}$ then implies that, in these normal coordinates and for radii $r$ of order $1/\sqrt{\mathsf{C}_{0}}$, at a given point $v\in\mathbb{B}_{\mathsf{g}}(p;r)$, the matrix $(\mathsf{g}_{ij}(v))_{i,j=1}^{\texttt{k}}$ representing the Riemannian metric remains uniformly close to the Euclidean metric $(\delta_{ij}(v))_{i,j=1}^{\texttt{k}}$ [^15] Here $\delta_{ij}$ denotes the Kronecker delta: $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise.. Consequently, the Riemannian volume element $\sqrt{\det(\mathsf{g}_{\mathcal{S}}(v))}\,dv$ is uniformly comparable from below to the Euclidean volume element $dv$ on such balls. Integrating this comparison over $\mathbb{B}_{\texttt{k}}(0;r)$ yields the stated lower bound on the Riemannian volume $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$. We now state the following key result.
@@
$|F_{\max}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)|\leq\frac{4L_{f,1}\delta^{\frac{1}{\texttt{k}}}}{(\kappa\cdot\mathsf{c}_{L}\cdot C_{\texttt{k}})^{\frac{1}{\texttt{k}}}}.$
Here $\mathsf{c}_{L}$ and $c_{1}$ are defined in Section˜4.1, $\kappa$ in Section˜3.2, k is the dimension of $\mathcal{S}(\theta)$, $C_{\texttt{k}}=\pi^{\texttt{k}/2}/\Gamma(\texttt{k}/2+1)$, and $r_{0},\mathsf{C}_{0}$ are from Section˜3.2.
@@ 5 Oracle complexity of optimizing $F_{\max}$
To evaluate $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$, we use its variational representation in (2), which involves solving the inner minimization problem $\min_{\beta\in\mathbb{R}}\phi_{\lambda,\delta}(\theta,\beta)$. Since $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex but possibly non-smooth, a natural choice is to apply the subgradient method. [^18] Because the problem is one-dimensional, the bisection method is also applicable, provided the domain is bounded. To ensure convergence, the non-smooth optimization must be restricted to a bounded interval. In Section˜5.1, we show that the optimal set $\operatorname*{arg\,min}_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ lies within $\mathrm{B},\mathrm{B}]$, so that
@@
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz continuous, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the classical convergence guarantees for convex Lipschitz objectives (nemirovskij1983problem; lan2012validation) together with the large-deviation analysis for stochastic approximation (juditsky2008large, Thm. 2.1). Accordingly, the output $\hat{\beta}$ of projected stochastic gradient descent (PSGD) (see Algorithm˜1) satisfies
@@
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\hat{\epsilon}^{-2}\log(\Delta^{-1}))$ iterations.
@@ 5.1 Inner minimization of Gibbs-superquantile
Suppose that Section˜2 holds. Given $\theta\in\Theta$, by using Algorithm˜1 ($\mathrm{PSGD}(\theta,\beta_{0};(\lambda,\delta))$), with the number of iterations $L=\mathcal{O}\left(\texttt{k}\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log\left(\epsilon_{v}^{-1}\right)\right)$, and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2-2\texttt{k}}$, we obtain
@@
5.2 Projected Zeroth-order algorithm with minima selection
In this subsection, we analyze the projected stochastic zeroth-order method with minima selection (PSZO-MinSel) introduced in Algorithm 2. Our goal is to quantify how many oracle calls are needed for PSZO-MinSel to reach an approximate Goldstein stationary point of $F_{\max}$.
Algorithm 2 Projected Stochastic Zeroth-order with minima selection (PSZO-MinSel)
@@ 5.2 Projected Zeroth-order algorithm with minima selection
where $\tilde{\mathsf{g}}_{\rho}(\theta_{n}):=b_{u}^{-1}\sum_{t=1}^{b_{u}}\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}),$ and $\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}):=\frac{m}{2\rho}\big[\tilde{\psi}(\theta_{n}+\rho u_{n,t})-\tilde{\psi}(\theta_{n}-\rho u_{n,t})\big]\,u_{n,t}$. Here $\tilde{\psi}$ is defined in (24), and $\{u_{n,t}\}_{t=1}^{b_{u}}$ are i.i.d. draws from the uniform distribution on the unit sphere $\mathbb{S}^{m-1}$. The complete procedure is summarized in Algorithm 2. Concretely, lines 4–5 invoke PSGD to approximate the minimisers of the perturbed sub-problems $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}\!+\!\rho u,\beta)$ and $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}-\rho u,\beta)$. These two inner solutions are combined in line 8 to form the standard two-point stochastic gradient estimator of $\tilde{F}_{{\texttt{SQ-G}}}$, which then drives the projected descent update for $\theta$ in line 9.
@@
Let $\{\theta_{i}\}_{i=1}^{N}$ be the sequence generated by Algorithm˜2. Under Sections˜2, 2 and 2, we obtain $\min_{i\in[1:N]}\mathbb{E}\left[\min_{g\in\partial_{\rho}F_{\max}(\theta_{i})}\|\mathcal{G}_{\Theta}(\theta_{i},g;\eta)\|\right]=\mathcal{O}(\epsilon),$ (27) with $N\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}\geq 1$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$, and $L\geq\mathcal{O}\left(\frac{\texttt{k}m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\cdot\log\left(m(\rho\epsilon)^{-1}\right)\right).$
@@
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{3+2\texttt{k}}\cdot(\texttt{k}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+4}\rho^{2\texttt{k}+4}}\right).$ (28)
@@ 5.3 Gradient Oracle Complexity under LMC-based Implementation
In the previous subsections, we analyze the sampling oracle complexity to identify an $(\epsilon,\rho,\eta)$-GGSP. To have a more concrete complexity result, we analyze the lower-level gradient oracle complexity when using Langevin Monte Carlo (LMC) to implement the Gibbs sampling oracle: To sample from the Gibbs measure in Equation˜3, a common choice is the Langevin dynamics
@@
where $W(t)$ denotes a $d$-dimensional Brownian motion. Note that the Gibbs measure in Equation˜3 is the unique invariant distribution of the above dynamics under mild conditions (chiang1987diffusion). In practice, one uses the Euler-Maruyama discretization of the above dynamics, which is also known as the unadjusted LMC method:
@@
Our idea is to replace the ideal Gibbs oracle in Algorithms˜1 and 2 by the output of LMC algorithm. While the analyses of this more concrete implementation closely follow the ones in the previous two subsections, we further need to ensure two points:
@@
    When $n$ sufficiently large, $\hat{\mu}_{n}$ is close to $\mu_{\theta}^{\lambda}$ in Wasserstein-1 distance;
@@
    Conditioned on the first point, make sure that replacing Gibbs samples with LMC samples introduces only a controlled bias in the Gibbs–superquantile estimator.
To fulfill the first point, we first use a general inequality relating $\mathbb{W}_{1}$, $\mathbb{W}_{2}$, and the Rényi divergence. By liu2020poincare,
@@
(chewi2024logconcave, Theorem 6.2.9) Consider the LMC iteration
@@
For the update (30), we pick $G(x):=g(\theta,x)/\lambda$ and $h_{k}=h/\lambda$. Then one has $\mathbb{R}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})\!\leq\!\varepsilon^{2}$ with $n=\tilde{\mathcal{O}}(dC_{PI}^{2}\lambda^{-2}\varepsilon^{-2})$. Combining this with the preceding Wasserstein bound shows that $\mathbb{W}_{1}(\hat{\mu}_{n},\mu^{\lambda}_{\theta})\leq 2C_{PI}^{\frac{1}{2}}C^{\frac{1}{2}}\varepsilon,$ where $C=2(e^{1/2}-1)$. Moreover, as discussed in the end of Section˜2.1.1, the inverse Poincaré constant $C_{PI}$ in our context is $\mathcal{O}(1)$, independent of $\lambda$ (gong2024poincare).
To establish the second point, we show in Appendix˜E that replacing the ideal Gibbs oracle in Algorithms 1 and 2 by the output of LMC algorithm after $n$ steps only introduces an additional approximation error proportional to $\delta^{-1}\mathbb{W}_{1}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})$. Choosing $n$ so that this error is of the same order as the statistical error controlled in Section˜5.1 ensures that all bounds in ˜5.5 remain valid, up to constants, when using LMC instead of exact Gibbs samples. This yields a concrete lower-level gradient oracle complexity, summarized in the following theorem.
@@
with $n\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}=\mathcal{O}(1)$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$,
$L=\mathcal{O}\left(\frac{m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\log(m\epsilon^{-1}\rho^{-1})\right),$
and the number of LMC steps $N=\tilde{\mathcal{O}}(dm^{6+6\texttt{k}}(\epsilon\rho)^{-6-6\texttt{k}})$. Moreover, the total number queries from lower-level gradient oracle is
$n\times N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{dm^{9+8\texttt{k}}[\texttt{k}\mathrm{B}^{2}+\tilde{C}_{f}]\log(\epsilon^{-1})}{\epsilon^{8\texttt{k}+10}\rho^{8\texttt{k}+10}}\right).$
The theorem above shows that Algorithm˜2 preserves its intrinsic-dimension complexity guarantees when the ideal Gibbs oracle is replaced by an LMC-based implementation, while making explicit the additional polynomial dependence on the ambient dimension $d$ and the factors $m^{\texttt{k}}$ and $(\rho\epsilon)^{-\texttt{k}}$.

### v4

@@ 3.2 Common Dimension and Geometric Properties of $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$
For each $\theta\in\Theta$, the local PŁ condition implies that on $\mathcal{S}(\theta)$ the Hessian $\partial_{x}^{2}g(\theta,x)$ is positive definite in the normal directions and degenerate in the tangent directions [73, Proposition 2.7]. In particular, its rank $\texttt{r}(\theta)$ is constant over $x\in\mathcal{S}(\theta)$, and the dimension of the minimizer manifold can be written as $\dim\mathcal{S}(\theta)=\texttt{k}(\theta)=d-\texttt{r}(\theta)$. Moreover, the solution mapping $\theta\mapsto\mathcal{S}(\theta)$ is continuous (Section˜3.1), and $\partial_{x}^{2}g(\theta,x)$ is continuous in $(\theta,x)$. Thus, when $\theta$ is perturbed slightly, points in $\mathcal{S}(\theta)$ move continuously and the corresponding Hessians remain close, so no eigenvalue can cross zero and the rank $\texttt{r}(\theta)$ cannot change. Hence $\texttt{r}(\theta)$, and therefore $\texttt{k}(\theta)$, is locally constant in $\theta$ and we obtain a single intrinsic dimension shared by all manifolds $\{\mathcal{S}(\theta)\}_{\theta\in\Theta}$. We summarize this in the following lemma.
@@ 4.1 Approximation of $F_{\max}$ with $\tilde{F}_{\texttt{SQ}}$
Classical results such as [36] give a pointwise asymptotic expansion of the volume of a small geodesic ball, showing that $\mathcal{M}(B_{\mathsf{g}}(p;r))/r^{\texttt{k}}\to c(p)\,$ as $r\to 0$ with a coefficient depending on the local curvature at $p$. For our purposes, this is not sufficient: we will see that in the proof of ˜4.2 we need a uniform, non-asymptotic lower bound that holds simultaneously for all centers $p\in\mathcal{S}(\theta)$ and all radii $0<r\leq r_{1}$, with $r_{1}$ depending only on the global bounds on sectional curvature and injectivity radius from Section˜3.2. Section˜4.1 provides exactly this quantitative version by turning the asymptotic comparison into an explicit estimate with controlled radius and constants.
@@
The proof of Section˜4.1 is given in Section˜C.1. Based on Section˜4.1, the Riemannian volume of a geodesic ball with radius $r$, $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$, scales as $r^{\texttt{k}}$. At a high level, the lemma relies on two geometric ingredients: (i) The injectivity radius bound $r_{0}>0$ guarantees that, around every $p\in\mathcal{S}$, we can introduce normal coordinates on the whole geodesic ball $\mathbb{B}_{\mathsf{g}}(p;r)$ for all $r\leq r_{0}$. (ii) The curvature bound $|\mathrm{Sec}|\leq\mathsf{C}_{0}$ then implies that, in these normal coordinates and for radii $r$ of order $1/\sqrt{\mathsf{C}_{0}}$, at a given point $v\in\mathbb{B}_{\mathsf{g}}(p;r)$, the matrix $(\mathsf{g}_{ij}(v))_{i,j=1}^{\texttt{k}}$ representing the Riemannian metric remains uniformly close to the Euclidean metric $(\delta_{ij}(v))_{i,j=1}^{\texttt{k}}$ [^15] Here $\delta_{ij}$ denotes the Kronecker delta: $\delta_{ij}=1$ if $i=j$ and $\delta_{ij}=0$ otherwise.. Consequently, the Riemannian volume element $\sqrt{\det(\mathsf{g}_{\mathcal{S}}(v))}\,dv$ is uniformly comparable from below and above to the Euclidean volume element $dv$ on such balls. Integrating this comparison over $\mathbb{B}_{\texttt{k}}(0;r)$ yields the stated bounds on the Riemannian volume $\mathcal{M}\bigl(\mathbb{B}_{\mathsf{g}}(p;r)\bigr)$. We now state the following key result.
@@
$|F_{\max}(\theta)-\tilde{F}_{\texttt{SQ}}(\theta)|\leq\frac{4L_{f,1}\delta^{\frac{1}{\texttt{k}}}}{(\kappa\cdot\mathsf{c}_{L}\cdot C_{\texttt{k}})^{\frac{1}{\texttt{k}}}},$
where $\mathsf{c}_{L}$ and $c_{1}$ are defined in Section˜4.1, $\kappa$ in Section˜3.2, k is the dimension of $\mathcal{S}(\theta)$, $C_{\texttt{k}}=\pi^{\texttt{k}/2}/\Gamma(\texttt{k}/2+1)$, and $r_{0},\mathsf{C}_{0}$ are from Section˜3.2.
@@ 5 Oracle Complexity of Optimizing $F_{\max}$
To evaluate $\tilde{F}_{{\texttt{SQ-G}}}(\theta)$, we use its variational representation in (2), which involves solving the inner minimization problem $\min_{\beta\in\mathbb{R}}\phi_{\lambda,\delta}(\theta,\beta)$. Since $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex but possibly nonsmooth, a natural choice is to apply the subgradient method [^18] Because the problem is one-dimensional, the bisection method is also applicable, provided the domain is bounded.. To ensure convergence, the nonsmooth optimization must be restricted to a bounded interval. In Section˜5.1, we show that the optimal set $\operatorname*{arg\,min}_{\beta}\phi_{\lambda,\delta}(\theta,\beta)$ lies within $\mathrm{B},\mathrm{B}]$, so that
@@
Given that (i) $\phi_{\lambda,\delta}(\theta,\cdot)$ is convex and Lipschitz continuous, (ii) the feasible set is bounded, and (iii) the stochastic subgradient $\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)$ satisfies $|\tilde{\partial}_{\beta}\phi_{\lambda,\delta}(\theta,\beta;X)|\leq\sigma:=1+\delta^{-1}$, we invoke the classical convergence guarantees for convex Lipschitz objectives [51, Proposition 2]. Accordingly, the output $\hat{\beta}$ of projected stochastic gradient descent (PSGD) (see Algorithm˜1) satisfies
@@
after at most $\mathcal{O}(\mathrm{B}^{2}\sigma^{2}\hat{\epsilon}^{-2}\log^{2}(\Delta^{-1}))$ iterations.
@@ 5.1 Inner Minimization of Gibbs-superquantile
Suppose that Section˜2 holds. Given $\theta\in\Theta$, by using Algorithm˜1 ($\mathrm{PSGD}(\theta,\beta_{0};(\lambda,\delta))$), with the number of iterations $L=\mathcal{O}\left(\texttt{k}^{2}\mathrm{B}^{2}\epsilon_{v}^{-2-2\texttt{k}}\log^{2}\left(\epsilon_{v}^{-1}\right)\right)$, and $|I|\geq 4\tilde{C}_{f}\epsilon_{v}^{-2-2\texttt{k}}$, we obtain
@@
5.2 Projected Zeroth-order Algorithm with Minima-Selection
In this subsection, we analyze the projected stochastic zeroth-order method with minima-selection (PSZO-MinSel) introduced in Algorithm 2. Our goal is to quantify how many oracle calls are needed for PSZO-MinSel to reach an approximate Goldstein stationary point of $F_{\max}$.
Algorithm 2 Projected Stochastic Zeroth-order with Minima-selection (PSZO-MinSel)
@@ 5.2 Projected Zeroth-order Algorithm with Minima-Selection
where $\tilde{\mathsf{g}}_{\rho}(\theta_{n}):=b_{u}^{-1}\sum_{t=1}^{b_{u}}\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}),$ and $\tilde{\mathsf{g}}_{\rho}(\theta_{n},u_{n,t}):=\frac{m}{2\rho}\big[\tilde{\psi}(\theta_{n}+\rho u_{n,t})-\tilde{\psi}(\theta_{n}-\rho u_{n,t})\big]\,u_{n,t}$. Here $\tilde{\psi}$ is defined in (24), and $\{u_{n,t}\}_{t=1}^{b_{u}}$ are i.i.d. draws from the uniform distribution on the unit sphere $\mathbb{S}^{m-1}$. The complete procedure is summarized in Algorithm 2. Concretely, lines 4–5 invoke PSGD to approximate the minimisers of the perturbed sub-problems $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}\!+\!\rho u,\beta)$ and $\min_{\beta}\phi_{\lambda,\delta}(\theta_{n}-\rho u,\beta)$ [^19] If $\tilde{\psi}$ is defined only on $\Theta$, then the perturbed evaluations $\tilde{\psi}(\theta_{n}\pm\rho u_{n,t})$ can be ill-posed when $\theta_{n}\pm\rho u_{n,t}\notin\Theta$. A standard workaround is to use the interiorized feasible set $\Theta_{\rho}:=\{\theta\in\Theta:\theta+\rho\mathbb{B}_{m}(0;1)\subseteq\Theta\}$ and replace $\theta_{n}$ in the oracle queries by $\bar{\theta}_{n}:=\text{Proj}_{\Theta_{\rho}}(\theta_{n})$, so that $\bar{\theta}_{n}\pm\rho u_{n,t}\in\Theta$.. These two inner solutions are combined in line 8 to form the standard two-point stochastic gradient estimator of $\tilde{F}_{{\texttt{SQ-G}}}$, which then drives the projected descent update for $\theta$ in line 9.
@@
Let $\{\theta_{i}\}_{i=1}^{N}$ be the sequence generated by Algorithm˜2. Under Sections˜2, 2 and 2, we obtain $\min_{i\in[1:N]}\mathbb{E}\left[\min_{g\in\partial_{\rho}F_{\max}(\theta_{i})}\|\mathcal{G}_{\Theta}(\theta_{i},g;\eta)\|\right]=\mathcal{O}(\epsilon),$ (27) with $N\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}\geq\mathcal{O}(m^{2}\epsilon^{-2})$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$, and $L\geq\mathcal{O}\left(\frac{\texttt{k}^{2}m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\cdot\log^{2}\left(m(\rho\epsilon)^{-1}\right)\right).$
@@
$N\times(b_{u}L+b_{u}|I|)=\mathcal{O}\left(\frac{m^{5+2\texttt{k}}\cdot(\texttt{k}^{2}\mathrm{B}^{2}+\tilde{C}_{f})\cdot\log^{2}\left(m\epsilon^{-1}\rho^{-1}\right)}{\epsilon^{2\texttt{k}+6}\rho^{2\texttt{k}+4}}\right).$ (28)
@@ 5.3 Gradient Oracle Complexity under LMC-based Implementation
In the previous subsections, we analyze the sampling oracle complexity to identify an $(\epsilon,\rho,\eta)$-GGSP. For a more concrete complexity result, we analyze the lower-level gradient oracle complexity when using Langevin Monte Carlo (LMC) to implement the Gibbs sampling oracle: To sample from the Gibbs measure in Equation˜3, a common choice is the Langevin dynamics
@@
where $W(t)$ denotes a $d$-dimensional Brownian motion. Note that the Gibbs measure in Equation˜3 is the unique invariant distribution of the above dynamics under mild conditions [17]. In practice, one uses the Euler-Maruyama discretization of the above dynamics, which is also known as the unadjusted LMC method:
@@
Our idea is to replace the ideal Gibbs oracle in Algorithms˜1 and 2 by the output of LMC algorithm. While the analyses of this more concrete implementation closely follow the ones in the previous two subsections, we further need to ensure the following two points hold:
@@
    For a sufficiently large $n$, $\hat{\mu}_{n}$ is close to $\mu_{\theta}^{\lambda}$ in Wasserstein-1 distance;
@@
    Conditioned on the first point, replacing Gibbs samples with LMC samples introduces only a controlled bias in the Gibbs–superquantile estimator.
To fulfill the first point, we first use a general inequality relating $\mathbb{W}_{1}$, $\mathbb{W}_{2}$, and the Rényi divergence. By [63, Theorem 1.1],
@@
[16, Theorem 6.2.9] Consider the LMC iteration
@@
For the update (30), we pick $G(x):=g(\theta,x)/\lambda$ and $h_{k}=h/\lambda$. Then one has $\mathbb{R}_{2}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})\!\leq\!\varepsilon^{2}$ with $n=\tilde{\mathcal{O}}(dC_{PI}^{2}\lambda^{-2}\varepsilon^{-2})$. Combining this with the preceding Wasserstein bound shows that $\mathbb{W}_{1}(\hat{\mu}_{n},\mu^{\lambda}_{\theta})\leq 2C_{PI}^{\frac{1}{2}}C^{\frac{1}{2}}\varepsilon,$ where $C=2(e^{1/2}-1)$. Moreover, as discussed in the end of Section˜2.1.1, the inverse Poincaré constant $C_{PI}$ in our context is $\mathcal{O}(1)$, independent of $\lambda$ [35].
To establish the second point, we show in Appendix˜E that the additional bias resulting from replacing the ideal Gibbs oracle in Algorithms 1 and 2 by the output of an $n$-step LMC run is bounded by $\delta^{-1}\mathbb{W}_{1}(\hat{\mu}_{n},\mu_{\theta}^{\lambda})$. We then choose $n$ large enough so that this LMC-induced error is at most a constant multiple of the accuracy $\epsilon_{v}$ used in Section˜5.1. With this choice, the total oracle error remains of the same order as in the ideal Gibbs-sampling analysis, so the convergence guarantee of ˜5.5 carries over to the LMC-based implementation with only constant-factor changes. This yields a concrete lower-level gradient oracle complexity, summarized in the following theorem.
@@
with $N\geq\mathcal{O}(m\epsilon^{-2}\rho^{-2})$, $b_{u}\geq\mathcal{O}(m^{2}\epsilon^{-2})$, $|I|\geq\tilde{C}_{f}m^{2\texttt{k}+2}(\rho\epsilon)^{-2\texttt{k}-2}$, $\eta\leq\mathcal{O}(m^{-1/2}\rho)$,
$L=\mathcal{O}\left(\frac{\texttt{k}^{2}m^{2\texttt{k}+2}\mathrm{B}^{2}}{(\rho\epsilon)^{2\texttt{k}+2}}\log^{2}(m\epsilon^{-1}\rho^{-1})\right),$
and the number of LMC steps $n=\tilde{\mathcal{O}}(dm^{6+6\texttt{k}}(\epsilon\rho)^{-6-6\texttt{k}})$. Moreover, the total number queries to the lower-level gradient oracle and the upper-level function oracle are
$N\times n\times(b_{u}L+b_{u}|I|)$ $=\mathcal{O}\!\left(\frac{d\,m^{11+8\texttt{k}}\,[\texttt{k}^{2}\,\mathrm{B}^{2}+\tilde{C}_{f}]\log^{2}(m\epsilon^{-1}\rho^{-1})}{\epsilon^{8\texttt{k}+12}\rho^{8\texttt{k}+10}}\right),$
$N\times b_{u}\times|I|$ $=\mathcal{O}\left(\frac{m^{2\texttt{k}+5}}{\rho^{2\texttt{k}+4}\epsilon^{2\texttt{k}+6}}\right),$
respectively.
The LMC layer only changes the cost of each oracle call. The theorem above shows that Algorithm˜2 preserves its intrinsic-dimension complexity guarantees when the ideal Gibbs oracle is replaced by an LMC-based implementation, while making explicit the additional polynomial dependence on the ambient dimension $d$ and the factors $m^{\texttt{k}}$ and $(\rho\epsilon)^{-\texttt{k}}$.